the value of a new car purchased for 36000 decreases 7.5% each year how long will it take for the value of the car to reach 20000
Answer by stanbon(75887)

the value of a new car purchased for 36000 decreases 7.5% each year how long will it take for the value of the car to reach 20000
------
Each year it retains 0.925 of its last-year value.
-------
Ans: V(t) = 36000*(0.925)^t
-----
Solve: 36000(0.925)^t = 20000
-----
t*log(0.925) = log(20/36)
----
t = log(5/9)/log(0.925)
-----
t = 7.54 years
=================
Cheers,
Stan H.